OpenStudy (anonymous):

The results from spinning a spinner multiple times are given in the table below. Color Frequency blue 7 yellow 14 green 13 red 9 purple 17 What is the experimental probability of not spinning purple?

OpenStudy (anonymous):

The sample space has 60 outcomes ( 7 + 14 + 13 + 9 + 17 ). Purple occurred 17 /60 . Therfore P ( NOT Purple ) = 1 - 17/60 = 43/60